OpenStudy (lena772):

Y=1.154x+0.769 rearrange to resolve for y = x + y, my x is -0.666

OpenStudy (rhystic):

Are you asking to solve for y & x? because resolving for y = x + y utilizing y=1.154+.769 is impossible unless x = 0. If you are just solving for x,y using y = 1.154x+.769 then youre slope (m) is correct.

OpenStudy (triciaal):

please verify the original question

OpenStudy (lena772):

I'm solving for X=-0.666+y, but Idk the y value @rhystic @triciaal

TheSmartOne (thesmartone):

Are you trying to find the solution to: Y=1.154x+0.769 X=-0.666+y

OpenStudy (lena772):

ues

OpenStudy (lena772):

i need to kow the value for y

TheSmartOne (thesmartone):

So, which do you prefer... substitution, elimination, or better yet, solving by graphing?

OpenStudy (lena772):

I have this graphed

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/f2k3enwgmr

TheSmartOne (thesmartone):

On the graph, it gives you the point of intersection.

OpenStudy (lena772):

When I put 0y it was wrong though

TheSmartOne (thesmartone):

It isn't 0 The point of intersection is (-0.6688, -0.0028) And so what is the y-value of the intersection?

OpenStudy (lena772):

-0.0028?

TheSmartOne (thesmartone):

correct :)

OpenStudy (lena772):

So do you add those together to find the answer or is that the answer

TheSmartOne (thesmartone):

that is the answer :)